OpenStudy (anonymous):

Find the point on the line 3 x + 3 y + 4 =0 which is closest to the point ( -3, 5 ).

OpenStudy (anonymous):

We have to find the orthogonal line, how can we find this line? \[3x+3y+4=0\] \[y=-x-\frac{4}{3}~~~~~~(Line~1)\] \[y-y_o=m*(x-x_o)\] \[m=1\] \[y-5=x-(-3)\] \[y=x+8~~~~~~(Line~2)\] now we have to find where this two lines will be the same value \[-x-\frac{4}{3}=x+8\] \[-\frac{4}{3}-8=2x\] \[\boxed{\boxed{x=-\frac{14}{3}}}\] now we have to find the Y, to do this verification we can use both equation, I'll choose the first one \[y=-x-\frac{4}{3}\] \[y=-\left(-\frac{14}{3}\right)-\frac{4}{3}\] \[y=\frac{14}{3}-\frac{4}{3}\] \[\boxed{\boxed{y=\frac{10}{3}}}\] \[\therefore~P=\left(-\frac{14}{3},\frac{10}{3}\right)\]